# InspectorRAGet: An Introspection Platform for RAG Evaluation

### v1

@@ 5 Conclusion
We present InspectorRAGet, a publicly available444https://github.com/IBM/InspectorRAGet platform to empower researchers, developers and stakeholders to gain a deeper understanding of the strengths and limitations of RAG systems. It includes aggregate-level and instance-level views as well as capabilities to explore human and algorithmic metrics and annotator behavior, allowing for more holistic analysis. We describe two pertinent use cases - RAG model evaluation and LLM-as-a-Judge - that showcase and highlight the value of these features, and publicly release our use-case input files and notebooks. In the future, we will explore adding additional capabilities to InspectorRAGet to facilitate comprehensive pattern discovery, as well as extending analytical capabilities beyond RAG to other popular LLM tasks such as summarization and code generation.
@@ 6 Ethical Considerations
A key claim in our paper is to include human annotations in the evaluation process. Human evaluation is subjective and prone to errors. We expect even the best annotators to make mistakes. We address mitigating this by including multiple annotators per question, but biases still may occur.
@@ Appendix A Experiment Results File Format
As InspectorRAGet is a web application, we naturally gravitated towards adopting JSON as an input format. Our prescribed structure for an experiment results file is intuitive and strives to minimize repetition of information.
@@
The experiment result file can be broadly split into six sections along the functional boundaries. The first section captures general details about the experiment in name, description and timestamp fields. The second and third sections describe the sets of models and metrics used in the experiment via the models and metrics fields, respectively. The last three sections cover the dataset and the outcome of evaluation experiment in the form of documents, tasks and evaluations fields. Figure 3 describes the requirements associated with each section in detail.
Appendix B Sample Experiment Runner
To help users employ InspectorRAGet as part of the broader RAG evaluation life cycle, we will also release a sample experiment runner in the form of a Python notebook. The notebook demonstrates how to run experiments using Hugging Face 555https://huggingface.co assets (datasets, models, and metric evaluators) and output the results in the experiment result JSON format expected by the platform. For our example, we use the HuggingFaceH4/ultrafeedback_binarized dataset, run inference using the zephyr-7b-sft-lora and Mistral-7B-v0.1 models, and calculate Bert score, RougeL, and BLEU metrics. The notebook can be easily modified by changing the dataset (or picking one from Hugging Face datasets), adding other models for comparison, and/or adding more metrics. Once the model predictions and metrics scores are generated, we convert them to the format outlined in Appendix A and generate a file that can be imported into InspectorRAGet.
Appendix C InspectorRAGet Implementation
InspectorRAGet is a React666https://react.dev web application built with NextJS 13 framework777https://nextjs.org. We extensively use the Carbon Design System888https://carbondesignsystem.com, an open-source design system with a wide range of assets including react and web components, styling guidelines, custom icons, and others. Adopting such a design system grants us the freedom to concentrate our efforts on user experience (UX) over user interface (UI).
As mentioned in Section 2, since InspectorRAGet, does not run the evaluation experiment (relying instead on the experiment results being provided to it as input), it does not require GPUs to run. This allows the platform to be deployed on lightweight virtual machines or even on personal laptops/desktops.
Finally, InspectorRAGet is built as a stateless application and does not retain any uploaded datasets. This choice was made with privacy in mind. Employing a stateless approach allows organizations to deploy a single hosted instance of the platform for multiple teams/members without the risk of sensitive data or evaluation results being inadvertently shared to inappropriate parties.
Appendix D Use Case Screenshots

### v2

@@ 5 Conclusion
We present InspectorRAGet, a publicly available888https://github.com/IBM/InspectorRAGet platform to empower researchers, developers, and stakeholders to gain a deeper understanding of the strengths and limitations of RAG systems. It includes aggregate-level and instance-level views as well as capabilities to explore human and algorithmic metrics and annotator behavior, allowing for a more holistic analysis. Our evaluation shows InspectorRAGet’s ability to yield concrete and actionable insights on two pertinent use cases - analyzing RAG model performance and LLM-as-a-judge performance. We publicly release our platform, input files used in this paper, and notebooks. In the future, we will explore augmenting InspectorRAGet to facilitate comprehensive pattern discovery, as well as extending its analytical capabilities beyond RAG to other popular LLM tasks, such as summarization and code generation.
@@ 6 Ethical Considerations
A key claim in our paper is to include human annotations in the evaluation process. Human evaluation is subjective and prone to errors. We expect even the best annotators to make mistakes. We suggest mitigating this by including multiple annotators per question, but biases still may occur.
@@ Appendix A Experiment Results File Format
As InspectorRAGet is a web application, we naturally gravitated towards adopting JSON as the input format. Our prescribed structure for an experiment results file is intuitive and strives to minimize repetition of information.
@@
The experiment result file can be broadly split into six sections along their functional boundaries. The first section captures general details about the experiment, including its name, description, and timestamp. The second and third sections describe the sets of models and metrics used in the experiment, respectively. The fourth and fifth sections cover the dataset, in the form of a list of documents/passages that are used in the experiment and a list of tasks, each representing a data instance, which is composed of the user input (e.g., question/conversation), references to the corresponding documents/passages, and, optionally, a list of reference responses. Finally, the sixth section includes information about the outcome of the evaluation, in the form of the scores of the different evaluation metrics for each task.
Note, that as part of the experiment results file, one is not providing the implementation of the metrics and models used, but rather high-level metadata about them (e.g., the name of a metric and its scale - e.g., yes/no, numeric - and the name of a model) along with their outputs (i.e., the resulting evaluation scores and model responses). This separation of the evaluation experiment implementation and run from the analysis of the results, allows InspectorRAGet to be agnostic of the specific models or metrics used, thus allowing the analysis of diverse evaluation experiments.
A detailed description of the input format can be found on our GitHub repository 999https://github.com/IBM/InspectorRAGet.
Appendix B Sample Experiment Runners
To help users employ InspectorRAGet as part of the broader RAG evaluation life cycle, we have also released sample experiment runners, showcasing how to use InspectorRAGet in combination with popular evaluation frameworks. Each experiment runner is provided in the form of a Python notebook, which demonstrates how to use the corresponding evaluation framework to run an evaluation experiment and transform its output to the input format expected by InspectorRAGet for an analysis of the evaluation results. As of this writing, we have released notebooks demonstrating integrations of our platform with the following popular frameworks:
Language Model Evaluation Harness 101010https://github.com/EleutherAI/lm-evaluation-harness Gao et al. (2023); a popular evaluation framework used to evaluate LLMs on different tasks.
RAGAs 111111https://github.com/explodinggradients/ragas Es et al. (2023); a popular evaluation framework specifically designed for the evaluation of RAG systems through LLM-as-a-judge techniques.
HuggingFace 121212https://huggingface.co Wolf et al. (2020), which offers libraries and assets (incl. datasets, models, and metric evaluators) that can be used to both create and evaluate RAG systems.
All experiment runners are available on the platform’s GitHub repository.